OpenStudy (anonymous):

WILL MEDAL! Which expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = Choose exactly two answers that are correct. A. 14x – 15 B. 2(7x – 10) C. 15x – 14 D. 2(7x – 9)

OpenStudy (anonymous):

you get no solution when you get some false statement like 0 = 1 . so if you simplify the left side you get : 24x - 18 - 10x 14x - 18 so now look at the answers: A 14x - 15 you get : -18 = - 15 no solution. B 14x - 20 you get -18 = -20 no solution C 15x-14 you get : 14x - 18 = 15x - 14 x = -4 there is a solution D 14x - 18 so 0 = 0 there is a solution (in fact infinite solutions)

OpenStudy (anonymous):

So the answer will be C and D

OpenStudy (anonymous):

I mean A or B

OpenStudy (anonymous):

A and B :)

OpenStudy (anonymous):

Thank you :)